2x-1=y , 2y-4x=7
equations parallel to each other?

Answer by algebrahouse.com(1659)

Parallel lines have equal slopes.

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

2x - y = 1
-y = -2x + 1 {subtracted 2x from each side}
y = 2x - 1 {divided each side by -1}
slope is 2

2y - 4x = 7
2y = 4x + 7 {added 4x to each side}
y = 2x + (7/2) {divided each side by 2}
slope is 2

Slopes are equal. The lines are parallel.

For more help from me, visit: www.algebrahouse.com